# LEAP-VO: Long-term Effective Any Point Tracking for Visual Odometry

### v2

@@ 4.2 Implementation details
We employ CoTracker [18], a recent PIPs variant, as our base TAP model. We train our model on the TAP-Vid-Kubric [9] training set, following [18]. We load the pre-trained CoTracker weights for the image feature extractor and train LEAP for 100,000 steps using 4 NVIDIA A100 GPUs in parallel. During training, we use $K=4$ steps for iterative refinements, $N=256$ for the number of queries, $S=8$ for the model window, and $S_{LP}=12$ for the tracking window. For inter-track anchor-based attention, the number of anchors is 64. The loss weights are set to $w_{1}=1.0,w_{2}=0.5,w_{3}=0.5$. In LEAP-VO, we perform 4 steps for each bundle adjustment with the Huber Loss for distance metric. The bundle adjustment window size $S_{BA}$ is set to 15. Track filtering parameters are set at $\gamma_{v}=0.9,\gamma_{d}=0.9,\gamma_{u}=0.8,\gamma_{track}=3$. Please refer to our supplementary materials for more details.
@@ 4.3 Visual Odometry Results
We compare our methods against the state-of-the-art SLAM and VO approaches, including ORB-SLAM2 [24], DynaSLAM [1], DROID-SLAM [32], TartanVO [40], DytanVO [28], and DPVO [33]. Evaluations are conducted on the static Replica dataset [29] as well as on the dynamic datasets MPI-Sintel [4] and TartanAir-Shibuya [26].
Replica. We compare our method with other baselines on the Replica dataset, which contains eight static scenes without any moving objects. Challenges arise in two test sequences that contain prolonged occlusions, leading to the potential camera tracking loss. As illustrated in Table 1, our approach outperforms other VO methods on all three metrics, showing effectiveness under static environments and the capability to handle occlusions. Further insights are available in Section 4.5. For classical methods, ORB-SLAM2 [24] and DynaSLAM [1], we observe tracking loss and resuming via re-localization, which reduces long-term drift and shows high accuracy for camera pose tracking.
MPI-Sintel. We also evaluate our VO camera trajectory accuracy on 14 dynamic scenes from the MPI Sintel dataset, following the setting of [48]. As shown in Table 2, our method substantially surpasses other VO baselines in dynamic scenes, thanks to our temporal uncertainty estimation and dynamic track detection. Traditional feature-based methods like ORB-SLAM2 [24] and DynaSLAM [1] sometimes fail to track camera movements in highly dynamic sequences, primarily due to outliers from dynamic object correspondences and segmentation-based only dynamic track estimation. In contrast, our approach, which considers visual, temporal, and inter-track information, demonstrates notable improvements in dynamic scene tracking accuracy.
TartanAir-Shibuya. Following DytanVO [28], we further evaluate the VO performance using the dynamic TartanAir-Shibuya dataset. We assess the Average Trajectory Error (ATE) across seven diverse scenes, ranging from simple to challenging. As depicted in Table 3, our method consistently outperforms all other SLAM and VO baselines in the majority of these scenes. Moreover, trajectory visualizations in Figure 5 also highlight the robust camera tracking capabilities of our method in dynamic environments.
@@ 4.4 Dynamic Track Estimation Results
Figure 6: Visualization of dynamic track estimation on DAVIS [19], MPI-Sintel [4], and TartanAir-Shibuya [26]. Odd columns: all point trajectories. Even columns: estimated dynamic point trajectories.
@@ 4.5 Ablation Study
Table 4: Effect of the proposed track filtering method on MPI-Sintel [4]. Tracking using all criteria yields the best results.
@@
Policy: Uniform Random SIFT [22] Ours
@@
Table 5: Comparison of keypoint extraction policies ($N_{a}=64$) on TartanAir-Shibuya [26]. Our distributed image-gradient sampling method achieves better results compared to other policies.
Baseline comparison. We evaluate the pretrained CoTracker [18] for the end VO performance. Table 6 shows that LEAP front-end outperforms CoTracker on all datasets, demonstrating the effectiveness of the proposed dynamic track estimation and temporal probability modeling.
@@
Table 6: Baseline comparison with CoTracker front-end [18]. Our LEAP front-end achieves better results on all datasets.

### v3

@@ 4.2 Implementation details
We employ CoTracker [karaev2023cotracker], a recent PIPs variant, as our base TAP model. We train our model on the TAP-Vid-Kubric [doersch2022tap] training set, following [karaev2023cotracker]. We load the pre-trained CoTracker weights for the image feature extractor and train LEAP for 100,000 steps using 4 NVIDIA A100 GPUs in parallel. During training, we use $K=4$ steps for iterative refinements, $N=256$ for the number of queries, $S=8$ for the model window, and $S_{LP}=12$ for the tracking window. For inter-track anchor-based attention, the number of anchors is 64. The loss weights are set to $w_{1}=1.0,w_{2}=0.5,w_{3}=0.5$. In LEAP-VO, we perform 4 steps for each bundle adjustment with the Huber Loss for distance metric. The bundle adjustment window size $S_{BA}$ is set to 15. Track filtering parameters are set at $\gamma_{v}=0.9,\gamma_{d}=0.9,\gamma_{u}=0.8,\gamma_{track}=3$. Please refer to our supplementary materials for more details.
@@ 4.3 Visual Odometry Results
We compare our methods against the state-of-the-art SLAM and VO approaches, including ORB-SLAM2 [murorbslam2], DynaSLAM [bescos2018dynaslam], DROID-SLAM [teed2021droid], TartanVO [wang2021tartanvo], DytanVO [shen2023dytanvo], and DPVO [teed2022deep]. Evaluations are conducted on the static Replica dataset [straub2019replica] as well as on the dynamic datasets MPI-Sintel [Butler2012sintel] and TartanAir-Shibuya [qiu2022airdos].
Replica. We compare our method with other baselines on the Replica dataset, which contains eight static scenes without any moving objects. Challenges arise in two test sequences that contain prolonged occlusions, leading to the potential camera tracking loss. As illustrated in Table 1, our approach outperforms other VO methods on all three metrics, showing effectiveness under static environments and the capability to handle occlusions. Further insights are available in Section 4.5. For classical methods, ORB-SLAM2 [murorbslam2] and DynaSLAM [bescos2018dynaslam], we observe tracking loss and resuming via re-localization, which reduces long-term drift and shows high accuracy for camera pose tracking.
MPI-Sintel. We also evaluate our VO camera trajectory accuracy on 14 dynamic scenes from the MPI Sintel dataset, following the setting of [zhao2022particlesfm]. As shown in Table 2, our method substantially surpasses other VO baselines in dynamic scenes, thanks to our temporal uncertainty estimation and dynamic track detection. Traditional feature-based methods like ORB-SLAM2 [murorbslam2] and DynaSLAM [bescos2018dynaslam] sometimes fail to track camera movements in highly dynamic sequences, primarily due to outliers from dynamic object correspondences and segmentation-based only dynamic track estimation. In contrast, our approach, which considers visual, temporal, and inter-track information, demonstrates notable improvements in dynamic scene tracking accuracy.
TartanAir-Shibuya. Following DytanVO [shen2023dytanvo], we further evaluate the VO performance using the dynamic TartanAir-Shibuya dataset. We assess the Average Trajectory Error (ATE) across seven diverse scenes, ranging from simple to challenging. As depicted in Table 3, our method consistently outperforms all other SLAM and VO baselines in the majority of these scenes. Moreover, trajectory visualizations in Figure 5 also highlight the robust camera tracking capabilities of our method in dynamic environments.
@@ 4.4 Dynamic Track Estimation Results
Figure 6: Visualization of dynamic track estimation on DAVIS [KhoRohrSch_davis], MPI-Sintel [Butler2012sintel], and TartanAir-Shibuya [qiu2022airdos]. Odd columns: all point trajectories. Even columns: estimated dynamic point trajectories.
@@ 4.5 Ablation Study
Table 4: Effect of the proposed track filtering method on MPI-Sintel [Butler2012sintel]. Tracking using all criteria yields the best results.
@@
Policy: Uniform Random SIFT [lowe1999sift] Ours
@@
Table 5: Comparison of keypoint extraction policies ($N_{a}=64$) on TartanAir-Shibuya [qiu2022airdos]. Our distributed image-gradient sampling method achieves better results compared to other policies.
Baseline comparison. We evaluate the pretrained CoTracker [karaev2023cotracker] for the end VO performance. Table 6 shows that LEAP front-end outperforms CoTracker on all datasets, demonstrating the effectiveness of the proposed dynamic track estimation and temporal probability modeling.
@@
Table 6: Baseline comparison with CoTracker front-end [karaev2023cotracker]. Our LEAP front-end achieves better results on all datasets.